# Tool-as-Interface: Learning Robot Policies from Human Tool Usage through Imitation Learning

### v1

@@ IV Our Robot Policy Learning Framework
Our framework enables the direct transfer of human play data into deployable robot policies. It is designed to fulfill the following key objectives:
@@
  • Support for Dynamic and High-Precision Tasks: Human play, with its inherent fluidity, enables the execution of highly dynamic tasks. Examples include flipping an egg in a pan or performing other actions that require swift, accurate, and natural motions — challenges that are often difficult to address with traditional teleoperation systems or handheld grippers.
@@
The following sections outline the framework’s design (Figure 2), its underlying principles, and logistical considerations for development and deployment.
IV-A Tool Usage for Data Collection and Manipulation
Humans naturally and intuitively use tools for everyday tasks such as cooking, eating, cleaning, and interacting with the world. Tools act as extensions of human actions, enabling diverse interaction with objects. The natural relationship between tools and objects provides an ideal interface for training robots to mimic human actions using tools, with minimal gap between human and robot tool usage. While grasping and pick-and-place tasks have been extensively studied in previous works, our work focuses on enabling robots to use the same tools humans commonly employ to interact with their environment effectively with the following benefits:
  • Minimized Embodiment Gap: Abstracting actions to the tool pose reduces morphological dependency, enabling policies to generalize across embodiments.
  • Scalable Data Collection: Simplifies the data collection process by eliminating the need for costly robot-specific demonstrations, making our method more accessible.
During data collection, humans can naturally use tools with their hands without requiring additional devices. For robot deployment, the tool can be attached to the robot in two ways:
  • Rigid Grasping: Grasping or picking, which has been extensively studied in prior works, is demonstrated in our Kinova Gen3 robot experiments and involves the robot securely grasping the tool.
  • Customized Fast Tool Changer: Designed for versatility, the tool changer is compatible with any robot that uses the ISO 9409-1-50-4-M6 flange, as demonstrated in our UR5e experiments.
IV-B Embodiment-Agnostic Perception
To encourage cross-embodiment transfer, we adopt a strategy that reduces the perception gap between the training and deployment phases. During training, human play data is collected, featuring human hands interacting with tools and objects. In deployment, robots execute the learned tasks. As showcased in our experiments, the visual differences between human hands and robotic end-effectors can introduce discrepancies that hinder generalization. To address this, we employ Grounded-SAM [49] to segment and mask out the embodiments in each phase. During training, human hands are masked, while during deployment, the robotic embodiments are masked, which ensures that the remaining parts of the scene in both training and testing phases appear visually similar. By aligning perception across embodiments, the framework mitigates distractions caused by embodiment-specific features, enabling better generalization to human-to-robot policy transfer.
IV-C View Augmentation
We use cameras for data collection due to their availability. With approximately 7.14 billion smartphones equipped with cameras, our approach can scale effectively [30]. However, using data from a single camera introduces challenges such as a lack of 3D perception and sensitivity to camera pose.
3D Reconstruction
To address these issues, we use MASt3R [34], an image-matching model that reconstructs accurate 3D environments from two RGB images. This eliminates the need for additional depth sensors, which are less common and consume more power compared to RGB sensors. We use two cameras, which is the minimum requirement for 3D reconstruction, to capture demonstration data. Using only one camera can lead to scale ambiguity in monocular settings, where poses can be scaled by an arbitrary, scene-dependent factor. By capturing images from two viewpoints, MASt3R reconstructs a 3D point cloud without requiring camera extrinsics or intrinsics, and globally aligns point maps within a multi-view 3D reconstruction framework. The process results in high-quality 3D representations.
Data Augmentation
Using 3D Gaussian splatting, we model the scene and synthesize novel viewpoints from human interaction data, which effectively augments the dataset, generating additional perspectives even if the training data was captured from only two views. These synthetic viewpoints provide the robot with a multi-angle understanding of the scene, allowing the policy to be trained on a more diverse and comprehensive set of visual inputs. Additionally, we apply random cropping to the images for data augmentation before feeding them into the policy network, following the approach from diffusion policy [14, 15]. Random cropping further improves the method’s robustness, enabling the policy to generalize better to variations in visual inputs.
IV-D Action Representation for Tool Manipulation
To support general tool usage, we propose a task-frame, tool-centric action representation denoted as $T_{\text{tool}}^{\text{task}}$. This representation focuses on the tool being manipulated, independent of human or robot morphology or camera pose. A visual depiction of the coordinate systems is provided in Figure 3.
Using a 6D pose estimation model (e.g., FoundationPose [71]), we determine the tool’s pose in the camera frame, $T_{\text{tool}}^{\text{camera}}$. To make the policy robust to camera and base movement, we transform this into the task frame:
where $T_{\text{camera}}^{\text{task}}$ represents the transformation from the camera to the task frame.
@@ IV-D Action Representation for Tool Manipulation
TABLE I: Benchmark Attributes of Real-World Tasks. These benchmarks evaluate the precision, adaptability, and capability of our framework to address tasks requiring high precision, handling extreme dynamics, utilizing extrinsic dexterity, performing in contact-rich scenarios, and overcoming gravity.
@@
IV-E Training and Deploying Robot Policies
We use diffusion policy [14] as our policy representation to predict $T_{\text{tool}}^{\text{task}}$, trained using the ACCESS system [7]. During deployment, for stationary robots, the task frame aligns with the base frame. For robots with moving base, base movement is compensated using $T_{\text{task}}^{\text{base}}$. The final end-effector pose in the robot base frame, used as the control command for the robot controller, is computed as:
where $T_{\text{eef}}^{\text{tool}}$ is the fixed transformation between the tool and the robot end-effector.
Figure 3: Coordinate System Diagram. This diagram illustrates the Camera frame, Tool frame, Task space frame, End-Effector (EEF) frame, and Base frame. We use $T_{\text{tool}}^{\text{task}}$ as the action representation for the visuomotor policy output.
V Policy Evaluations
Our experimental evaluations aim to assess the effectiveness of our framework for deploying robot policy learning across three key dimensions:
  • Capabilities: What skills can our framework enable robots to acquire, and how robust are the policies under challenging conditions like a moving camera or base?
  • Effectiveness: How well does our framework achieve its objectives outlined in Section IV? Can it support reliable and scalable policy learning for complex tasks while streamlining data collection for highly dynamic, contact-rich, or dexterous scenarios?
  • Policy Execution Efficiency: How efficiently do the trained policies execute tasks? Does our framework enable smoother and more natural motion trajectories? Can it achieve faster and more fluid task completion compared to baseline methods?
To evaluate our framework for learning from direct human play, we developed a set of real-world robotic task domains using two embodiments: Kinova Gen3 and UR5e. These tasks are designed to test various aspects of policy capabilities, efficiency and effectiveness. Table I summarizes the key characteristics of these tasks. Visual input for the policies is provided by two RealSense D415 cameras. We describe the tasks in detail, highlighting their challenges and the specific capabilities we aim to test.
@@ V Policy Evaluations
Nail Hammering: The task involves hammering a 3D-printed nail, requiring the robot to locate the nail, draw back the hammer, and strike the nail tip accurately. With a diameter of less than 15.5 mm, the nail tip demands high precision. Challenges include localizing the nail tip precisely and planning effective hammer trajectories. To evaluate generalization, the initial position of the nail is varied across different spatial configurations. We collected 180 seconds of data (40 human play episodes) from a single participant.
Meatball Scooping: In this task, the robot must use a spoon to scoop a meatball from a pan and transfer it to a bowl. This task is challenging due to the complex dynamics of the meatball, which can roll unpredictably within the pan. Additionally, the interaction between the spoon and the meatball requires careful control, as improper contact can cause the meatball to slip or escape the spoon. We randomize the initial position of the meatball within the pan to test its generalization capability. We collected 340 seconds of data (50 human play episodes) from a single participant.
Pan Flipping (Egg, Burger Bun, Meat Patty): The objective of this task is to use a pan to flip various objects, such as an egg, a burger bun, and a meat patty. The task is challenging due to its high-speed dynamics, requiring the robot to overcome gravity and accurately manage the interaction between the pan and the objects. Each object differs in weight, shape, and texture, adding further complexity. This task evaluates the policy’s ability to handle fast, contact-rich interactions and adapt to diverse object types. To increase variability, the initial positions of the objects within the pan are randomized. Furthermore, the rapid and dynamic nature of the task makes it unsuitable for classical demonstration collection methods, highlighting the advantages of using bare-handed human play for data collection. We collected 50 seconds of data (38 human play episodes) from a single participant using three different pans and two object types.
Wine Balancing: In this task, the robot needs to use a hook to lift a wine bottle and carefully insert it into an unstable, zero-gravity wine rack. The task is challenging due to the precise control required to suspend the bottle in mid-air and counteract gravitational forces effectively. Any over-insertion or under-insertion will cause the bottle to lose balance. To constrain the horizontal movement of the rack, screws were added as obstacles to limit lateral motion. No additional variability was introduced. We collected 223 seconds of data (15 human play episodes) from a single participant.
Soccer Ball Kicking: In this task, the robot must use a golf club to kick a ball that slides into a field and direct it into the goal. To increase the challenge, a 3D-printed row of players serves as obstacles between the robot and the goal. The task is difficult because the robot must accurately intercept the moving ball, strike it with the correct force and direction, and ensure it avoids obstacles before reaching the goal. The position of the player obstacle varies. We collected 78 seconds of data (20 human play episodes) from a single participant.
Baselines: The primary focus of this work is to demonstrate the effectiveness and efficiency of learning directly from human play without relying on robot-generated data. We compare our approach against two baselines: a diffusion policy trained on robot demonstrations, and UMI [16], a hand-held gripper-based data collection method. The robot demonstration dataset is collected using either a SpaceMouse or Gello interface under the same data collection time. Additionally, we perform ablation studies to analyze the impact of key components, such as random cropping of images before policy training, novel view synthesis-based data augmentation, and the effects of embodiment segmentation. To further illustrate the advantages of our approach, we compare trajectory rollouts for a meatball-scooping episode, highlighting how our method is more sample-efficient and less prone to distribution shifts by eliminating excessive waypoints.
TABLE II: Task Success Rates and Completion Times. Success rates are the number of successful trials out of total episodes, and average completion times are based on successful trials. “DP” refers to the diffusion policy trained on teleoperation data. “Not Feasible” tasks denote cases where teleoperation failed due to extreme dynamics, precision, or reactivity demands. Our method consistently achieves higher success rates and shorter completion times.
Task Method Success Rate Time (s)
Hammer Nailing DP 0/13 -
Ours 13/13 11.0
Meatball Scooping DP 5/12 42.0
Ours 10/12 12.4
Pan Flipping - Egg DP Not Feasible -
Ours 12/12 1.5
Pan Flipping - Burger Bun DP Not Feasible -
Ours 9/12 1.9
Pan Flipping - Meat Patty DP Not Feasible -
Ours 10/12 2.3
Wine Balancing DP Not Feasible -
Ours 8/10 30.9
Soccer Ball Kicking DP Not Feasible -
Ours 6/10 2.0
TABLE III: Task success rates comparing our method with the hand-held gripper-based method on Nail Hammering.
Method Demo Duration & Count Success Rate
UMI [16] $\sim$180 seconds (25 demos) 0/13
UMI $\sim$720 seconds (100 demos) 13/13
Ours $\sim$180 seconds (40 demos) 13/13
VI Experiment Results
In our experiments, we demonstrate that our framework is both effective and efficient for training robots with advanced capabilities. Furthermore, leveraging human play data enables robots to perform smoother movements and acquire skills that are challenging or even impossible to achieve with robot-generated data.
Figure 6: Policy Execution Trajectory Comparison. (a) Initial setup for meatball scooping. (b) Comparison of end-effector XY trajectories from our framework and a policy trained on robot-collected data. Our method generates smoother, more natural, and efficient trajectories with fewer waypoints, enabling fluid and precise motion. In contrast, the robot-collected policy produces jerky movements with excessive waypoints, which hinder task performance.
VI-A Capabilities and Effectiveness
Table II presents the results of our real-world robot tasks, showing that our framework consistently outperforms baseline methods by achieving significantly higher success rates across all evaluated scenarios. We further compare our method with a stronger hand-held gripper baseline, UMI [16], as shown in Table III. In our default setup, SLAM-based mapping failed due to low environmental texture. To address this limitation, we added a textured background to support reliable mapping for UMI. For the nail hammering task, we evaluated UMI using 25 demonstrations (matching our collection time) and 100 demonstrations (to assess its ideal performance). UMI failed all 13 trials with 25 demonstrations but succeeded in all 13 trials with 100. UMI could not be applied to the wine balancing task due to contact-induced tool displacement, nor to the pan flipping task due to tool inertial slippage. In the soccer kicking task, large and fast motions made it nearly impossible to localize the demonstration trajectory within the initial map.
In contrast to the baselines, our method demonstrates reliable performance across all tasks. As illustrated in Figure 4, our method excels in real-world policy rollouts by accurately detecting spatial locations in tasks such as nail hammering and meatball scooping. For pan flipping tasks, it performs the high-speed motions required to effectively flip eggs, burger buns, and meat patties. Additionally, it demonstrates precise control in lifting and inserting a wine bottle into its stand and reacts swiftly and accurately to kick a soccer ball in the soccer ball kicking task. The superior performance of our framework is primarily due to its ability to collect a significantly larger volume of episodes within the same data collection timeframe, as detailed later in Section VII. This broader dataset covers a wider range of task variations, enabling more robust and adaptable policy training. Furthermore, our approach overcomes the inherent limitations of Gello and SpaceMouse, enabling the collection of demonstrations for scenarios that these tools cannot adequately handle due to their constraints.
VI-B Policy Execution Efficiency
Our framework demonstrates exceptional efficiency in task execution, achieving faster task completion times and producing smoother action motions compared to baseline methods, as shown in Table II. The efficiency is largely attributed to the nature of human play data, which captures the fluidity and speed of real-world human activities, resulting in smoother and more natural trajectories in the training dataset. In contrast, previous approaches relying on teleoperated data often suffer from significantly slower speeds and less natural motions, limiting their effectiveness in dynamic scenarios. By leveraging the realistic dynamics of human play, our framework not only accelerates task execution but also enhances motion quality, making it better suited for real-world applications.
@@ VI-B Policy Execution Efficiency
Figure 7: Initial States for All Evaluation Episodes. All methods are evaluated using the same set of manually defined initial states, overlayed in the image. These states ensure diverse variations to test the policy’s spatial generalization capabilities.
Figure 9: Human Perturbation Robustness. This figure showcases the robot’s ability to handle human-induced perturbations across three tasks: (1) In nail hammering, the robot successfully followed a manually moved nail; (2) In meatball scooping, it located and scooped meatballs even when additional ones were thrown into the pan mid-task; and (3) In egg flipping, the robot consistently flipped the egg back after human intervention repositioned it. These results highlight the policy’s robustness to unpredictable human perturbations.
VI-C Benefits of Tool-Based Action Representation in Task Space
We observed that using the tool in the camera frame as the action representation allows the policy to perform comparably to prior works when the camera is stationary. However, the success rate drops to zero when the camera is moving. Tracking the camera pose in real time during motion is non-trivial, leading to incorrect calculations of the end-effector position in the base frame. When the base moves, policies relying on the tool in the base frame fail entirely, achieving zero success. These failures occur because such policies assume a fixed relative transformation between the base and the workspace, an assumption invalidated by base motion. In contrast, representing actions using the tool in task space enables the policy to remain effective under base perturbations. Our approach demonstrates robust performance when camera poses vary, when the base moves, and even during simultaneous shaking of the camera and base.
VI-D Effects of Random Cropping and View Augmentation
Our experiments show that random cropping (RC) and view augmentation (VA) enhance policy robustness to diverse camera poses. Training with RC improves resilience to minor camera perturbations, such as small movements or shaking, while VA extends robustness by exposing the model to a broader range of camera configurations during training. We evaluated these techniques on the nail hammering task and present the results in Figure 5, comparing three models: one trained with both RC and VA, one with RC only, and one without either technique. The combined use of RC and VA significantly expands the working range of camera configurations, including those with substantial deviations from the original viewpoint. The robustness arises from the model’s exposure to a diverse set of camera poses during training.
VI-E Policy Execution Trajectory Comparison
Our framework generates faster, smoother, and more natural trajectories compared to traditional approaches, as demonstrated by the end effector (EEF) XY trajectory for the meatball scooping task in Figure 6. Figure 6(a) depicts the initial setup of one trial, while Figure 6(b) contrasts the XY trajectory of the policy rollout using our framework with that of a policy trained on robot-collected data. Our framework surpasses the baseline approach in two critical aspects:
  • Improved Motion Fluidity: Trajectories generated by our framework are significantly smoother and less jerky, enabling more fluid and precise task execution. In contrast, the robot-collected data results in jerky motions caused by excessive waypoints, which disrupt movement precision and degrade performance.
  • Efficient and Effective Learning: The trajectories produced by our framework contain 10 times fewer waypoints than those from robot-collected data. The reduction simplifies the learning space, mitigates cumulative error growth, and minimizes the risk of distribution shifts commonly observed in behavior cloning. As a result, our policies exhibit significantly improved sample efficiency.
VI-F Generalization
Spatial Generalization: Our policy demonstrates strong generalization across various spatial configurations. Figure 7 illustrates the range of initial states tested for multiple tasks. In the hammer nailing task, nail positions vary. For meatball scooping, meatball positions are tested. In soccer ball kicking, we evaluate two goalkeeper configurations. For pan flipping, object poses vary to test all areas of the pan. As shown in Table II, our method achieves high success rates across these ranges.
Object Generalization: Our method generalizes effectively to different objects in pan flipping tasks, including a toy egg, a 3D-printed meat patty, and a real burger bun (Figure 7, second column). With only 13 demonstrations, the policy succeeds by leveraging a simple but effective strategy: tilting the pan to slide the object into a corner, then flicking the pan to propel and flip the object. Our manipulation approach for pan flipping enables robust generalization across diverse object types.
Tool Generalization: To assess the generalization ability of our policy across different tools, we conducted a pan-flipping experiment using a burger bun and five pans: large, medium, small, tiny, and square. For each pan, we collected 12 trials with varying initial configurations and reported the success rates (Fig. 10). The policy was trained on demonstrations using the large, medium, and square pans, and evaluated on all five. Results indicate that our method exhibits some generalization across both pan sizes and shapes (circular vs. square). High success rates were observed with the large and medium pans. However, performance declined on smaller pans, likely due to their limited surface area. The square pan also showed lower success rates, as its shallow edges allowed the bun to slide out during flipping.
@@ VI-F Generalization
Figure 10: Tool Generalization (a) The tested pans. (b) Success rate across 12 testing trials.

### v2

@@ A.1 Key Capabilities and Practical Benefits
Our framework enables the direct transfer of human manipulation data into deployable robot policies. It is designed to fulfill the following key objectives:
@@
  • Support for Dynamic and High-Precision Tasks: Human manipulation, with its inherent fluidity, enables the execution of highly dynamic tasks. Examples include flipping an egg in a pan or performing other actions that require swift, accurate, and natural motions — challenges that are often difficult to address with traditional teleoperation systems or handheld grippers.
@@
Appendix B Detailed Experiment Setup
B.1 Task Descriptions
Nail Hammering: The task involves hammering a 3D-printed nail, requiring the robot to locate the nail, draw back the hammer, and strike the nail tip accurately. With a diameter of less than 15.5 mm, the nail tip demands high precision. Challenges include localizing the nail tip precisely and planning effective hammer trajectories. To evaluate generalization, the initial position of the nail is varied across different spatial configurations. We collected 180 seconds of data (40 episodes) from a single participant.
Meatball Scooping: In this task, the robot must use a spoon to scoop a meatball from a pan and transfer it to a bowl. This task is challenging due to the complex dynamics of the meatball, which can roll unpredictably within the pan. Additionally, the interaction between the spoon and the meatball requires careful control, as improper contact can cause the meatball to slip or escape the spoon. We randomize the initial position of the meatball within the pan to test its generalization capability. We collected 340 seconds of data (50 episodes) from a single participant.
Pan Flipping (Egg, Burger Bun, Meat Patty): The objective of this task is to use a pan to flip various objects, such as an egg, a burger bun, and a meat patty. The task is challenging due to its high-speed dynamics, requiring the robot to overcome gravity and accurately manage the interaction between the pan and the objects. Each object differs in weight, shape, and texture, adding further complexity. This task evaluates the policy’s ability to handle fast, contact-rich interactions and adapt to diverse object types. To increase variability, the initial positions of the objects within the pan are randomized. Furthermore, the rapid and dynamic nature of the task makes it unsuitable for classical demonstration collection methods, highlighting the advantages of using bare-handed human videos for data collection. We collected 50 seconds of data (38 episodes) from a single participant using three different pans and two object types.
Wine Balancing: In this task, the robot needs to use a hook to lift a wine bottle and carefully insert it into an unstable, zero-gravity wine rack. The task is challenging due to the precise control required to suspend the bottle in mid-air and counteract gravitational forces effectively. Any over-insertion or under-insertion will cause the bottle to lose balance. To constrain the horizontal movement of the rack, screws were added as obstacles to limit lateral motion. No additional variability was introduced. We collected 223 seconds of data (15 episodes) from a single participant.
Soccer Ball Kicking: In this task, the robot must use a golf club to kick a ball that slides into a field and direct it into the goal. To increase the challenge, a 3D-printed row of players serves as obstacles between the robot and the goal. The task is difficult because the robot must accurately intercept the moving ball, strike it with the correct force and direction, and ensure it avoids obstacles before reaching the goal. The position of the player obstacle varies. We collected 78 seconds of data (20 episodes) from a single participant.
@@ B.1 Task Descriptions
Table 3: Benchmark Attributes of Real-World Tasks. These benchmarks evaluate the precision, adaptability, and capability of our framework to address tasks requiring high precision, handling extreme dynamics, utilizing extrinsic dexterity, performing in contact-rich scenarios, and overcoming gravity.
@@
Figure 11: Initial States for All Evaluation Episodes. All methods are evaluated using the same set of manually defined initial states, overlaid in the image. These states ensure diverse variations to test the policy’s spatial generalization capabilities.
B.2 Implementation Details
B.2.1 Hardware Design
B.2.2 Tool Pose Estimation
Figure 12: Policy Execution Trajectory Comparison. (a) Initial setup for meatball scooping. (b) Comparison of end-effector XY trajectories from our framework and a policy trained on robot-collected data.
@@ B.2.2 Tool Pose Estimation
Figure 13: Tool Generalization. (a) The tested pans. (b) Success rate across 12 testing trials.
Appendix C Additional Experimental Results
[…]
Effects of Embodiment Segmentation: Embodiment Segmentation masks the agent’s embodiments during data collection and policy deployment, ensuring visually consistent scenes and reducing the training-deployment visual gap. Embodiment Segmentation significantly improves policy performance, as shown in Figure 14. Figure 14(a) highlights failure cases without segmentation. In the wine balancing task, the robot strikes the table, triggering safety stops due to improper bottle handling. In the soccer ball kicking task, the robot’s actions are inconsistent, shorter, and less precise than during training. Quantitative results in Figure 14(b) further underscore segmentation’s impact. Across 10 trials, segmentation enabled 8 successes in the wine balancing task, while the model without it achieved none. Similarly, in the soccer ball kicking task, segmentation resulted in 6 successes, compared to 2 without it. By aligning training and testing visual distributions, Embodiment Segmentation ensures consistent and reliable robot performance during the training and deployment.
Appendix D Detailed Analysis on Data Collection Efficiency and Affordability
D.1 Data Collection Efficiency
Our framework achieves significantly higher data collection throughput than traditional methods, enabling more demonstrations within the same timeframe. The improvement is driven by the natural and intuitive efficiency of human manipulation, which ensures faster and more reliable task execution. Figure 7(a) highlights the superior manipulation capabilities of human hands, while Figure 8 quantifies the substantial time savings per episode. For nail hammering and meatball scooping, Gello and Spacemouse were used as teleoperation methods, respectively. Human hands reduced data collection time by 73% and 81% for nail hammering and meatball scooping, with consistently low variation in performance. In more complex tasks like pan flipping, wine balancing, and soccer ball kicking, teleoperation methods failed entirely due to limitations such as lack of tactile feedback, delays, and difficulty handling dynamic or precise actions. Our method further reduces data collection time by 41% compared to handheld grippers such as UMI [14] in nail hammering. UMI proved ineffective in wine balancing and pan flipping due to tool inertial slippage or contact-induced displacement, and failed in soccer kicking because of difficulty localizing large, fast motions. Moreover, it requires rich textures to build a pre-collection map, which our method does not. These results underscore the superior efficiency, robustness, and versatility of human manipulation as a scalable solution for high-quality robot learning datasets.
D.2 Reliability
Figure 7(b) and Figure 7(c) illustrates typical failure cases with Gello, Spacemouse, and UMI [14], which frequently encounter issues such as safety stops or collisions during data collection. In contrast, our method ensures smooth, uninterrupted operation, avoiding these limitations. Traditional methods face significant challenges in high-speed or complex tasks. For example, Gello and Spacemouse struggle with replicating the extreme dynamics and precise motions required for flipping objects like eggs during pan flipping, often resulting in unsuccessful attempts. Similarly, teleoperation delays prevent timely strikes during soccer ball kicking, consistently leading to missed kicks and repeated failures. In tasks like wine balancing, the absence of tactile feedback impairs precision during the data collection, causing the wine bottle to tip over during data collection. Furthermore, in meatball scooping, the velocity vectors generated by Spacemouse input lead to jerky trajectories with redundant waypoints, significantly reducing efficiency. These challenges make effective training impractical with traditional methods. By leveraging human manipulation, our framework not only addresses these limitations but also provides a reliable and scalable solution for dynamic and precision-demanding tasks.
D.3 Discussion of Data Collection Methods
Table 4 compares various data collection methods based on cost, usability, expertise requirements, intuitiveness, and precision. Our method incurs no additional cost (\$0), unlike hardware-dependent solutions like UMI and LEGATO, which demand significant investment. This affordability makes our approach accessible to users from diverse backgrounds without financial constraints. Unlike hardware-based systems such as UMI, LEGATO, Gello, and Spacemouse, which are prone to malfunctions and maintenance issues, our hardware-free framework ensures reliability and eliminates repair delays or expenses. Additionally, it requires no supplementary 3D printing, in contrast to approaches like Visual Imitation Made Easy, UMI, and LEGATO. The simplicity of our design promotes inclusivity in collecting large-scale dataset for robot learning research. Our method also offers a more natural experience compared to tools like Spacemouse, while being far more cost-effective than VR and AR devices. Moreover, systems like Gello and Spacemouse lack the precision necessary for dynamic tasks, a limitation addressed by our approach. Overall, our method is a cost-effective, and accessible solution for data collection, overcoming key drawbacks of existing approaches while reducing complexity and maintenance needs.